y''-8y'+20y=100x2-26xex solve the given differential equation by undetermined cofficient


1
Expert's answer
2021-11-30T06:14:21-0500

The corresponding homogeneous differential equation


"y''-8y'+20y=0"

The auxiliary equation


"r^2-8r+20=0"

"(r-4)^2=-4"

"r=4\pm2i"

The general solution of the corresponding homogeneous differential equation is


"y_h=c_1e^{4x}\cos(2x)+c_1e^{4x}\sin(2x)"

Find the particular solution of the nonhomogeneous differential equation


"y_p=Ax^2+Bx+C+(Dx+E)e^x"

"y_p'=2Ax+B+(Dx+E+D)e^x"

"y_p''=2A+(Dx+E+2D)e^x"

Substitute


"2A+(Dx+E+2D)e^x"

"-8(2Ax+B+(Dx+E+D)e^x)"

"+20(Ax^2+Bx+C+(Dx+E)e^x)"

"=100x^2-26xe^x"

"x^2:20A=100=>A=5"

"x^1:-16A+20B=0=>B=4"

"x^0:2A-8B+20C=0=>C=\dfrac{11}{10}"

"xe^x:D-8D+20D=-26=>D=-2"

"e^x:E+2D-8E-8D+20E=0=>E=-\dfrac{12}{13}"

Then


"y_p=5x^2+4x+\dfrac{11}{10}-2xe^x-\dfrac{12}{13}e^x"

The general solution of the given nonhomogeneous differential equation is


"y=c_1e^{4x}\cos(2x)+c_1e^{4x}\sin(2x)"


"+5x^2+4x+\dfrac{11}{10}-2xe^x-\dfrac{12}{13}e^x"